OpenStudy (anonymous):

whats 2 to the -23 power devided by seven

OpenStudy (mathmale):

You are expected to show your own efforts before asking for help.

OpenStudy (anonymous):

but its hard

OpenStudy (anonymous):

ok, so i know you have to multiply 2 by -23 first

OpenStudy (mathmale):

Absolutely not! 2^(-23) is exponentiation, not multiplication.

OpenStudy (anonymous):

@ddelph1 Break it up into pieces. First, complete the problem \[2^{-23}\]

OpenStudy (anonymous):

@mathmale is correct @ddelph1 , this is not a multiplication problem.

OpenStudy (anonymous):

ooooooopppssss

OpenStudy (anonymous):

my b

OpenStudy (anonymous):

@mathmale Sorry, I forgot. HOORAY!!!! MATHMALE IS ONLINE!!!!! lol

OpenStudy (anonymous):

Now, @ddelph1 Do you know how to solve equations with exponents? You know what, it might be better to write the equation in actual symbol and numbers, instead of words.

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

\[\frac{ 2^{-23} }{ 7 }\]

OpenStudy (mathmale):

ddelp1: You've asked the question. If you seriously want to arrive at an answer, start thinking and showing more work of your own.

OpenStudy (anonymous):

Use PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction According to PEMDAS, you have to solve 2^(-23) first, and then divide that answer by 7

OpenStudy (anonymous):

What is \[2^{-23}\]

OpenStudy (anonymous):

im trying. Calm down

OpenStudy (anonymous):

@ddelph1 Openstudy is not a place where you will just get the answers and leave. We are only here to guide you to your answers, not to give you the answers.

OpenStudy (anonymous):

Now, what is \[2^{-23}\]

OpenStudy (anonymous):

-8,388,608 i think

OpenStudy (anonymous):

How did you calculate that?

OpenStudy (anonymous):

ok so i think i did it right this time

OpenStudy (anonymous):

-16777216

OpenStudy (anonymous):

How did you calculate that? Show your work @ddelph1 .

OpenStudy (anonymous):

@mathmale Am I doing this correctly? Normally I just get help on OS, i don't generally give help

OpenStudy (anonymous):

2x-2 23 times

OpenStudy (anonymous):

I'm sorry, what?

OpenStudy (anonymous):

2 multiplied by -2 23 times

OpenStudy (anonymous):

Okay, so complete the problem.

OpenStudy (anonymous):

@mathmale It's actually been a while since I've done problems with negative exponents, what he is doing is correct, right?

rebeccaxhawaii (rebeccaxhawaii):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (mathmale):

2^23 would signify "2 used as a multiplicand 23 times." 2^2=2*2 =4, 2^3 = 2*2*2, and so on. "2 multiplied by -2 23times is not correct.

OpenStudy (anonymous):

I got it!!!!!!!!!!!!!!

OpenStudy (anonymous):

-2396745.142

OpenStudy (anonymous):

baboomn

rebeccaxhawaii (rebeccaxhawaii):

that isnt direct that isnt the answer @Needhelpstudying

rebeccaxhawaii (rebeccaxhawaii):

but whatever floats your boat

OpenStudy (anonymous):

I'm surprised no one has mentioned \[\large \sf a^{-n}=\frac{1}{a^n}\]

OpenStudy (mathmale):

2^(-23) is exponentiation, not multiplication. Isn't that an application of \[\large \sf a^{-n}=\frac{1}{a^n}?\]